# No Free Lunch from Random Feature Ensembles

### v1

@@ 2.2 Degrees of Freedom
Following notation similar to (Atanasov et al., 2024) and (Bach, 2023), we will write expressions in terms of the “degrees of freedom” defined as follows:
$\operatorname{Df}_{n}(\kappa)$ $\equiv\sum_{t}\frac{\eta_{t}^{n}}{(\eta_{t}+\kappa)^{n}},$ $\operatorname{tf}_{n}(\kappa)$ $\equiv\sum_{t}\frac{\bar{w}_{t}^{2}\eta_{t}^{n}}{(\eta_{t}+\kappa)^{n}},$ $n$$\in\mathbb{N}.$ (8)
Intuitively, $\operatorname{Df}_{n}(\kappa)$ can be understood as a measures of how many modes of the kernel eigenspectrum are above a threshold $\kappa$, with the sharpness of the measurement increasing with $n$. $\operatorname{tf}_{n}$ is a similar measure with each mode weighted by the corresponding component of the target function.
3 Omniscient Risk Estimates for Random Feature Ensembles
3.1 The Bias-Variance Decomposition of $E_{g}$
We first review the omniscient risk estimate $E_{g}^{1}$ (superscript indicates $K=1$) for a single RFRR model. We do not derive this well-known result here, but rather direct the reader to a wealth of derivations, including references (Atanasov et al., 2024; Canatar et al., 2021; Simon et al., 2023; Adlam & Pennington, 2020; Rocks & Mehta, 2021; Hastie et al., 2022; Zavatone-Veth et al., 2022). Translating the risk estimate into our selected notation, we may write:
$E_{g}^{1}\approx\frac{1}{1-\gamma_{1}}\left[-\rho\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2})+(1-\rho)\kappa_{2}\operatorname{tf}_{1}(\kappa_{2})+\sigma_{\epsilon}^{2}\right]$ (9)
@@ 3.1 The Bias-Variance Decomposition of $E_{g}$
Assuming concentration of the kernel eigenfunctions, Defilippis et al. et. al. show that a dimension-free multiplicative error bound of the form $|\mathcal{E}_{g}^{1}-E_{g}^{1}|\leq\tilde{\mathcal{O}}(N^{-1/2}+P^{-1/2})\cdot E_{g}^{1}$, where $\mathcal{E}_{g}^{1}$ is the “true” risk and $E_{g}^{1}$ given in eq. 9, holds with high probability over the input data and random weights. We find that eq. 9 provides an accurate estimate of risk at finite $N,P$. The error formula can further be decomposed using a bias-variance decomposition with respect to the particular realization $\bm{Z}$ of random features:
@@
$\operatorname{Bias}_{z}^{2}$ $\equiv\mathbb{E}_{\bm{x}\sim\mu_{\bm{x}}}\left[(\mathbb{E}_{\bm{Z}}\left[f(\bm{x})\right]-f_{*}(\bm{x}))^{2}\right]+\sigma_{\epsilon}^{2}$ (13)
$\operatorname{Var}_{z}$ $\equiv\mathbb{E}_{\bm{Z}}\mathbb{E}_{\bm{x}\sim\mu_{\bm{x}}}\left[(f(\bm{x})-\mathbb{E}_{\bm{Z}}\left[f(\bm{x})\right])^{2}\right]$ (14)
While the learned function $f$ in the equation above also depends on the particular realization of the dataset $\mathcal{D}$, we do not include this in the Bias-Variance decomposition because we are explicitly interested in the variance due to the realization of a finite set of random features. Furthermore, the Bias and Variance written in equations 13, 14 are expected to concentrate over $\mathcal{D}$ (Atanasov et al., 2024; Adlam & Pennington, 2020; Lin & Dobriban, 2021). Omniscient estimates for the Bias and Variance are given explicitly in (Simon et al., 2023):
$\operatorname{Bias}^{2}_{z}$ $=\frac{-\kappa_{2}^{2}}{1-\gamma_{2}}\operatorname{tf}_{1}^{\prime}(\kappa_{2})+\frac{\sigma_{\epsilon}^{2}}{1-\gamma_{2}}\,,$ $\operatorname{Var}_{z}$ $=E_{g}^{1}-\operatorname{Bias}^{2}_{z}\,,$ (15)
@@ 3.2 Ensembling Reduces Variance of the Learned Estimator
3.3 More is Better in Random Feature Ensembles
Feature-bagging is a long-standing practice in ensemble learning methods. In the case of random forests, for example, subsampling of data dimensions leads to improved performance, despite reducing the size of each decision tree (Breiman, 2001). Here, we find the opposite behavior for RFRR ensembles at optimal ridge.
Theorem 1.
@@ 3.3 More is Better in Random Feature Ensembles
In the special case $K=1$, this reduces to the “more is better” theorem for single models proven in (Simon et al., 2023). Proof of this theorem follows from the omniscient risk estimate 16, and is provided in Appendix B.
We demonstrate monotonicity with $P$ and $N$ in Fig. 1, where we plot $E_{g}^{K}$ as a function of both sample size $P$ and the network size $N$ in ensembles of $\operatorname{ReLU}$ random feature models applied to a binarized CIFAR-10 image classification task (see Appendix D.2). While error may increase with $P$ or $N$ at a particular ridge value $\lambda$, error decreases monotonically provided that the ridge $\lambda$ is tuned to its optimal value. Theoretical learning curves are calculated using eq. 16, with eigenvalues $\eta_{k}$ and target weights $\bar{w}_{k}$ determined by computing the NNGP kernel corresponding to the infinite-feature limit of the $\operatorname{ReLU}$ RF model (see Appendix D.3 for details). Numerically, we verify that error monotonicity with $P$ and $N$ holds at the level of a 00-$1$ loss on the predicted classes of held-out test examples for both score-averaging and majority-vote ensembling over the predictors (see fig. S3).
We compare ridge-optimized error across ensemble sizes $K$ in fig. S1, finding that increasing sample size $P$ and network size $N$ are usually more effective than ensembling over multiple networks in reducing predictor error, indicating that for the binarized CIFAR-10 RFRR classification task, bias is the dominant contribution to error. Similarly, ensembling over multiple networks gives meager improvements in performance relative to increasing network size in deep feature-learning ensembles Vyas et al. (2023).
Figure 1: “More is better” in random feature ensembles. We perform $\operatorname{ReLU}$ RFRR on a binarized CIFAR-10 classification task and compare the empirical test risk to the omniscient risk estimate (eq. 16). (A) We fix $N=256$ and vary both $P$ and $K$. Color corresponds to the regularization $\lambda$. Markers show numerical experiments and dotted lines theoretical predictions. Error is monotonically decreasing with $P$ provided that the regularization $\lambda$ is tuned to its optimal value. (B) Same as (A) except that $P=256$ is fixed and $K$, $N$ are varied. Markers and error bars show mean and standard deviation over $50$ trials.
@@ 4 No Free Lunch from Random Feature Ensembles
It is immediate from eq. 16 that increasing the size of an ensemble reduces the error. However, with a fixed memory capacity, a machine learning practitioner is faced with the decision of whether to train a single large model, or to train an ensemble of smaller models and average their predictions. Here, we prove a “no free lunch” theorem which says that, given a fixed total number of features $M$ divided evenly among $K$ random feature models, then the lowest possible risk will always be achieved by $K=1$, provided that the ridge is tuned to its optimal value. Furthermore, this ridge-optimized error increases monotonically with $K$.
Theorem 2.
@@
The proof follows a similar strategy to the proof of theorem 1, and is provided in appendix B. We test this prediction by performing ensembled $\operatorname{ReLU}$ RFRR on the binarized CIFAR-10 classification task in figure 2. We find that increasing $K$ while keeping the total number of features $M$ fixed always degrades the optimal test risk. The strength of this effect, however, depends on the size of the training set. For larger training sets ($P\gg N$), the width of each ensemble member becomes the constraining factor in each predictor’s ability to recover the target function. However, when $P\ll N$, the optimal loss is primarily determined by $P$, so that optimal error only begins increasing appreciably with $K$ once $N=M/K\lesssim P$ (fig. 2 B). We again find similar behavior of the 00-$1$ loss under score-average and majority-vote ensembling (see fig. S6). While the ridge-optimized error is always minimal for $K=1$, we notice in fig. 2C that near-optimal performance can be obtained with $K>1$ over a wider range of $\lambda$ values, suggesting that ensembling may offer improved robustness in situations where fine-tuning of the ridge parameter is not possible. Further robustness benefits have been reported for regression ensembles of heterogeneous size (Ruben & Pehlevan, 2023).
@@
Corollary 1.
Let $E_{g}^{K}(N)$ be the test risk of an ensemble of $K$ RFRR models each with $N$ features given by eq. 16. Suppose $K^{\prime}>K$ or $N^{\prime}<N$. It follows from Theorems 1 and 2 that
@@
We demonstrate this result on synthetic tasks with power-law structure and on the binarized CIFAR-10 classification task in fig. S2. For ensembles in the over-parameterized regime ($N\gg P$), this bound appears to be tight. To understand why, we expand the omniscient risk estimate $E_{g}^{K}$ (eq. 16) in a power series about $\lambda=1/N=0$ in Appendix C.2, finding:
$E_{g}^{K}=-\frac{P{\kappa_{2}^{*}}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2}^{*})}{P-\operatorname{Df}_{2}(\kappa_{2}^{*})}+\lambda F(\kappa_{2}^{*},P)+\frac{P\kappa_{2}^{*}\operatorname{tf}_{1}(\kappa_{2}^{*})}{KN}+\mathcal{O}(\lambda^{2},\lambda/N,1/N^{2})\,,$ (20)
@@
where $\kappa_{2}^{*}$ satisfies $\operatorname{Df}_{1}(\kappa_{2}^{*})=P$, and we have set $\sigma_{\epsilon}=0$. We see that, at leading order in $\lambda$, $1/N$, risk depends on the ensemble size $K$ and model size $N$ only through the total number of features $KN$. So, while Theorem 2 guarantees that there can be no benefit to dividing features into an ensemble of smaller predictors, we don’t expect this to appreciably harm performance as long as the ensemble remains in the overparameterized regime ($N\gg P$) and optimal or near-optimal performance can be achieved with a small ridge parameter.
Figure 2: No Free Lunch from Random Feature Ensembles. We perform kernel RF regression on a binarized CIFAR 10 classification task. (A) We vary $K$ and $N$ while keeping total parameter count $M=1024$ fixed. The sample size $P$ is indicated above each plot. (B) Error $E_{g}^{K}$ optimized over the ridge parameter $\lambda$ increases monotonically with $K$ provided the total parameter count $M$ is fixed. Dashed lines show theoretical prediction using eq. 16 and markers and error-bars show mean and standard deviation of the risk measured in numerical simulations across 10 trials. (C) We show error as a function of $\lambda$ for each $K$ value simulated and $P=8192$. Dashed lines show theoretical prediction using eq. 16 and shaded regions show standard deviation of risk measured in numerical simulations across 10 trials.
5 Width-Bottlenecked Scaling of Random Feature Ensembles

### v2

@@ 2.2 Degrees of Freedom
Using notation similar to (Atanasov et al., 2024) and (Bach, 2023), we will write expressions in terms of the “degrees of freedom” defined as follows:
$\operatorname{Df}_{n}(\kappa)$ $\equiv\sum_{t}\frac{\eta_{t}^{n}}{(\eta_{t}+\kappa)^{n}},$ $\operatorname{tf}_{n}(\kappa)$ $\equiv\sum_{t}\frac{\bar{w}_{t}^{2}\eta_{t}^{n}}{(\eta_{t}+\kappa)^{n}}\,$, (11)
where $n\in\mathbb{N}$. Intuitively, $\operatorname{Df}_{n}(\kappa)$ can be understood as a measure of how many modes of the kernel eigenspectrum are above a threshold $\kappa$, with the sharpness of the measurement increasing with $n$. $\operatorname{tf}_{n}$ is a similar measure with each mode weighted by the corresponding component of the target function.
3 Deterministic Equivalent Error for Random Feature Ensembles
The error $\mathcal{E}_{g}^{K}$ of eq. 10 is a random quantity, depending both on the particular realization of the dataset $\mathcal{D}$ and the realization of the random feature vectors $\{\bm{v}_{n}^{k}\}$. Many recent works have found, however, that for large values of $P$ and $N$ the error $\mathcal{E}_{g}^{K}$ will concentrate over these sources of randomness. The error $\mathcal{E}_{g}^{K}$ can then be estimated by a deterministic equivalent error $E_{g}^{K}$, which depends on the dataset $\mathcal{D}$ and random features $\{\bm{v}_{n}^{k}\}$ only through their sizes $P$ and $N$. We will write $\mathcal{E}_{g}^{K}\simeq E_{g}^{K}$, with the symbol $\simeq$ meaning that the value on the right side is the deterministic quantity approximating the stochastic quantity on the left in the limit of large $P$ and $N$. The exact slack in this approximation varies in the literature, and is discussed in Appendix B.
We first review the deterministic equivalent approximation to the error $\mathcal{E}_{g}^{1}$ of a single RFRR model (superscript indicates $K=1$). We do not derive this well-known result here, but rather direct the reader to a wealth of derivations, including references (Atanasov et al., 2024; Canatar et al., 2021; Simon et al., 2023; Adlam & Pennington, 2020; Rocks & Mehta, 2021; Hastie et al., 2022; Zavatone-Veth et al., 2022; Defilippis et al., 2024). Translating the risk estimate into our selected notation, we have that $\mathcal{E}_{g}^{1}\simeq E_{g}^{1}$ with:
$E_{g}^{1}=\frac{1}{1-\gamma_{1}}\left[-\rho\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2})+(1-\rho)\kappa_{2}\operatorname{tf}_{1}(\kappa_{2})+\sigma_{\epsilon}^{2}\right]$ (12)
@@ 3 Deterministic Equivalent Error for Random Feature Ensembles
We find that eq. 12 provides an accurate estimate of risk at finite $N,P$.
3.1 The Bias-Variance Decomposition of $E_{g}$
The error formula can further be decomposed using a bias-variance decomposition with respect to the particular realization $\bm{Z}$ of random features:
@@ 3.1 The Bias-Variance Decomposition of $E_{g}$
$\operatorname{Bias}_{z}^{2}$ $\equiv\mathbb{E}_{\bm{x}\sim\mu_{\bm{x}}}\left[\left(\mathbb{E}_{\bm{Z}}\left[\hat{f}(\bm{x})\right]-f_{*}(\bm{x})\right)^{2}\right]+\sigma_{\epsilon}^{2}$ (16)
$\operatorname{Var}_{z}$ $\equiv\mathbb{E}_{\bm{Z}}\mathbb{E}_{\bm{x}\sim\mu_{\bm{x}}}\left[\left(\hat{f}(\bm{x})-\mathbb{E}_{\bm{Z}}\left[\hat{f}(\bm{x})\right]\right)^{2}\right]$ (17)
While the learned function $f$ in the equation above also depends on the particular realization of the dataset $\mathcal{D}$, we do not include this in the Bias-Variance decomposition because we are interested in the variance due to the realization of a finite set of random features. Furthermore, the Bias and Variance written in equations 16, 17 are expected to concentrate over $\mathcal{D}$ (Atanasov et al., 2024; Adlam & Pennington, 2020; Lin & Dobriban, 2021). The deterministic equivalent formulas for the Bias and Variance are given explicitly in (Simon et al., 2023), and may be written as:
$\operatorname{Bias}^{2}_{z}$ $\simeq\frac{-\kappa_{2}^{2}}{1-\gamma_{2}}\operatorname{tf}_{1}^{\prime}(\kappa_{2})+\frac{\sigma_{\epsilon}^{2}}{1-\gamma_{2}}\,$, (18)
$\operatorname{Var}_{z}$ $\simeq E_{g}^{1}-\operatorname{Bias}^{2}_{z}\,$, (19)
@@ 3.2 Ensembling Reduces Variance of the Learned Estimator
4 More is Better in Random Feature Ensembles
A ubiquitous observation in the modern practice of machine learning is that larger models and datasets lead to better performance. However, ensemble learning offers an additional complication that can break this trend. In the case of random forests, for example, subsampling of data dimensions leads to improved performance, despite reducing the expressivity of each decision tree (Breiman, 2001). Similarly, deep feature-learning ensembles have been empirically observed to violate the “bigger is better” principle (see fig. 8 from Vyas et al. (2023)). Here, we show that no such violations can occur in RFRR ensembles.
Theorem 4.1.
@@ 4 More is Better in Random Feature Ensembles
Proof of this theorem follows from the omniscient risk estimate 20, and is provided in Appendix C. We make the following remarks:
Remark 4.2.
In the special case $K=K^{\prime}=1$, this reduces to the “more is better” theorem for single models proven in Simon et al. (2023), and thus extends the notion that larger models and datasets improve performance from single models to RFRR ensembles.
Remark 4.3.
Improved accuracy of the constituent models that form an ensemble does not necessarily lead to improved accuracy of the ensemble-averaged prediction, which suppresses the variance of the individual predictors. Proof of theorem 4.1 therefore requires a separate treatment of the bias and variance terms, unlike the $K=K^{\prime}=1$ case proven by Simon et al.
We demonstrate monotonicity with $P$ and $N$ in Fig. 1, where we plot $E_{g}^{K}$ as a function of both sample size $P$ and the network size $N$ in ensembles of $\operatorname{ReLU}$ random feature models applied to a binarized CIFAR-10 image classification task (see Appendix E.2). While error may increase with $P$ or $N$ at a particular ridge value $\lambda$, error decreases monotonically provided that the ridge $\lambda$ is tuned to its optimal value. Theoretical learning curves are calculated using eq. 20, with eigenvalues $\eta_{k}$ and target weights $\bar{w}_{k}$ determined by computing the NNGP kernel corresponding to the infinite-feature limit of the $\operatorname{ReLU}$ RF model (see Appendix E.3 for details). Numerically, we verify that error monotonicity with $P$ and $N$ holds at the level of a 0-11 loss on the predicted classes of held-out test examples for both score-averaging and majority-vote ensembling over the predictors (see fig. S3).
We compare ridge-optimized error across ensemble sizes $K$ in fig. S1, finding that increasing sample size $P$ and network size $N$ are usually more effective than ensembling over multiple networks in reducing predictor error, indicating that for the binarized CIFAR-10 RFRR classification task, bias is the dominant contribution to error. Similarly, ensembling over multiple networks gives meager improvements in performance relative to increasing network size in deep feature-learning ensembles (Vyas et al., 2023).
Figure 1: “More is better” in random feature ensembles. We perform $\operatorname{ReLU}$ RFRR on a binarized CIFAR-10 classification task and compare the empirical test risk to the omniscient risk estimate (eq. 20). (A) We fix $N=256$ and vary both $P$ and $K$. Color corresponds to the regularization $\lambda$. Markers show numerical experiments and dotted lines theoretical predictions. Error is monotonically decreasing with $P$ provided that the regularization $\lambda$ is tuned to its optimal value. (B) Same as (A) except that $P=256$ is fixed and $K$, NN are varied. Markers and error bars show mean and standard deviation over 5050 trials.
@@ 5 No Free Lunch from Random Feature Ensembles
It is immediate from eq. 20 that increasing the size of an ensemble reduces the error. However, training additional models incurs a significant cost. With a fixed memory capacity, the machine learning practitioner is faced with the decision of whether to train a single large model, or to train an ensemble of smaller models. Here, we prove a “no free lunch” theorem which says that, given a fixed total number of features $M$ divided evenly among $K$ random feature models, then the lowest possible risk will always be achieved by $K=1$, provided that the ridge is tuned to its optimal value. Furthermore, this ridge-optimized error increases monotonically with $K$.
Theorem 5.1.
@@
Several remarks are in order:
Remark 5.2.
In the limit where $M\to\infty$, this result implies that no ensemble of RFRR models can outperform the limiting (infinite-feature) KRR model with optimal ridge.
Remark 5.3.
In section 7, we comment on the distinction between this result and known equivalences between infinite ensembles and single predictors of infinite size (Patil & LeJeune, 2024; LeJeune et al., 2020; Dern et al., 2024)
Remark 5.4.
While ensembling is the canonical strategy for reducing the variance of a predictor, the proof of theorem 5.1 relies on the fact that when the ridge parameter is scaled to keep bias constant, increasing model size decreases variance of the learned predictor faster than adding additional predictors to the ensemble.
Proof of theorem 5.1 follows a similar strategy to the proof of theorem 4.1, and is provided in Appendix C. We test this prediction by performing ensembled $\operatorname{ReLU}$ RFRR on the binarized CIFAR-10 classification task in figure 2. We find that increasing $K$ while keeping the total number of features $M$ fixed always degrades the optimal test risk. The strength of this effect, however, depends on the size of the training set. For larger training sets ($P\gg N$), the width of each ensemble member becomes the constraining factor in each predictor’s ability to recover the target function. However, when $P\ll N$, the optimal loss is primarily determined by $P$, so that optimal error only begins increasing appreciably with $K$ once $N=M/K\lesssim P$ (fig. 2 B). We again find similar behavior of the 0-11 loss under score-average and majority-vote ensembling (see fig. S6). While the ridge-optimized error is always minimal for $K=1$, we notice in fig. 2C that near-optimal performance can be obtained with $K>1$ over a wider range of $\lambda$ values, suggesting that ensembling may offer improved robustness in situations where fine-tuning of the ridge parameter is not possible. Further robustness benefits have been reported for regression ensembles of heterogeneous size (Ruben & Pehlevan, 2023).
Figure 2: No Free Lunch from Random Feature Ensembles. We perform kernel RF regression on a binarized CIFAR 10 classification task. (A) We vary $K$ and NN while keeping total parameter count $M=1024$ fixed. The sample size $P$ is indicated above each plot. (B) Error $E_{g}^{K}$ optimized over the ridge parameter $\lambda$ increases monotonically with $K$ provided the total parameter count $M$ is fixed. Dashed lines show theoretical prediction using eq. 20 and markers and error-bars show mean and standard deviation of the risk measured in numerical simulations across 10 trials. (C) We show error as a function of $\lambda$ for each $K$ value simulated and $P=8192$. Dashed lines show theoretical prediction using eq. 20 and shaded regions show standard deviation of risk measured in numerical simulations across 10 trials.
@@
Corollary 5.5.
Let $E_{g}^{K}(N)$ be the test risk of an ensemble of $K$ RFRR models each with NN features given by eq. 20. Suppose $K^{\prime}>K$ or $N^{\prime}<N$. It follows from Theorems 4.1 and 5.1 that
@@
Remark 5.6.
We emphasize that this is a uni-directional statement. It is common for a single larger model to outperform an ensemble of smaller models with greater total size, particularly in the underparameterized regime. See, for example, fig. S1.B.
@@
We demonstrate this result on synthetic tasks with power-law structure and on the binarized CIFAR-10 classification task in fig. S2.
6 Near-Optimality Conditions for RFRR Ensembles
Our results indicate that ensembling is always a sub-optimal allocation of model parameters. Theorem 5.1 guarantees that no ensemble of random-feature models can outperform a single model with the same total size when the ridge parameter is optimally tuned. Furthermore, corollary 5.5 states that an ensemble of smaller models can only outperform an ensemble of larger models if the ensemble of smaller models has greater total size. These results suggest that, with a fixed memory capacity, it is always beneficial to scale a predictor by increasing the size of the network. However, ensembling remains an attractive option in practice because they allow fully parallelized training and inference. It is therefore important to ask when ensembles can attain near-optimal performance. In section 6.1, we investigate RFRR ensembles in the overparameterized regime, finding that at leading order in $\lambda$ and $1/N$, error depends only on the total feature count $M=KN$. To study the underparameterized regime, we derive scaling laws for RFRR ensembles under joint scaling of ensemble size $K$ and model size $N$. While the optimal scaling law is always achieved by scaling model size with a fixed ensemble size, our analysis provides conditions on the data and task eigen-structure under which joint scaling of model size and ensemble size can achieve a near-optimal scaling law for the error.
6.1 Near-Optimal Risk in Overparameterized RFRR Ensembles
For ensembles in the overparameterized regime ($N\gg P$), the bound of corollary 5.5 appears to be tight, meaning that an ensemble of $K$ models of size $N$ performs similarly to a single model of size $M=NK$. To understand why, we expand the deterministic equivalent error formula $E_{g}^{K}$ (eq. 20) as power series in $1/N\gtrsim 0$ and $\lambda\approx 0$ in Appendix D.2, finding:
$E_{g}^{K}$ $=-\frac{P{\kappa_{2}^{*}}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2}^{*})}{P-\operatorname{Df}_{2}(\kappa_{2}^{*})}+\lambda F(\kappa_{2}^{*},P)+\frac{P\kappa_{2}^{*}\operatorname{tf}_{1}(\kappa_{2}^{*})}{KN}$
$+\mathcal{O}(\lambda^{2},\lambda/N,1/N^{2})\,$, (24)
@@ 6.1 Near-Optimal Risk in Overparameterized RFRR Ensembles
where $\kappa_{2}^{*}$ satisfies $\operatorname{Df}_{1}(\kappa_{2}^{*})=P$, and we have set $\sigma_{\epsilon}=0$. We see that, at leading order in $\lambda$, and $1/N$, risk depends on the ensemble size $K$ and model size $N$ only through the total number of features $KN$. So, while Theorem 5.1 guarantees that there can be no benefit to dividing features into an ensemble of smaller predictors, we don’t expect this to appreciably harm performance as long as the ensemble remains in the overparameterized regime ($N\gg P$) and optimal or near-optimal performance can be achieved with a small ridge parameter. While we do not provide exact conditions under which near-optimal performance is achieved at small ridge parameter here, our numerical experiments verify that error depends on $N$ and $K$ only through their product $M=NK$ (see agreement between dotted lines and green lines in fig. S2). Furthermore, Simon et al. argue that in the overparameterized limit $N\to\infty$, optimal ridge approaches zero for tasks with power-law structure under realistic conditions.
6.2 Scaling Laws for Underparameterized RFRR Ensembles
Figure 3: Graphical Depiction of how ensemble size $K$ and model size NN scale with total feature count $M$ under different growth exponents $\ell$ (see eq. 25).